# NSNQuant: A Double Normalization Approach for Calibration-Free Low-Bit Vector Quantization of KV Cache

### v1

@@ Abstract
Large Language Model (LLM) inference is typically memory-intensive, especially when processing large batch sizes and long sequences, due to the large size of key-value (KV) cache. Vector Quantization (VQ) is recently adopted to alleviate this issue, but we find that the existing approach is susceptible to distribution shift due to its reliance on calibration datasets. To address this limitation, we introduce NSNQuant, a calibration-free Vector Quantization (VQ) technique designed for low-bit compression of the KV cache. By applying a three-step transformation—1) a token-wise normalization (Normalize), 2) a channel-wise centering (Shift), and 3) a second token-wise normalization (Normalize)—with Hadamard transform, NSNQuant effectively aligns the token distribution with the standard normal distribution. This alignment enables robust, calibration-free vector quantization using a single reusable codebook. Extensive experiments show that NSNQuant consistently outperforms prior methods in both 1-bit and 2-bit settings, offering strong generalization and up to 3$\times$ throughput gain over full-precision baselines.
@@ 1 Introduction
Large language models (LLMs) have been widely adopted across various domains due to their strong generalization capabilities (achiam2023gpt, ). Recently, with the emerging trend of using LLMs to solve complex problems, their use has expanded into long-context scenarios such as long-context reasoning (wei2022chain, ; guo2025deepseek, ) and retrieval-augmented generation (RAG) (lewis2020retrieval, ). However, when processing long sequences, LLM inference requires a significant amount of memory and is memory-bound (kim2023squeezellm, ; hooper2024kvquant, ).
One of the major causes is the large size of key-value (KV) cache, which linearly increases with the sequence length. To alleviate this issue, many studies have proposed approaches to compressing KV cache effectively, such as eviction (zhang2023h2o, ; li2024snapkv, ; cai2024pyramidkv, ; qin2025cake, ) and low-rank approximation (changpalu, ; lin2024matryoshkakv, ). Among them, quantization is one of the most widely adopted approaches. Previous studies (liu2024kivi, ; hooper2024kvquant, ) analyze how outliers emerge in the KV cache and propose to quantize key and value along the channel dimension and token dimension, respectively. Recently, Coupled Quantization (CQ) (zhang2024kv, ) proposed to quantize multiple channels together, using centroids obtained from the calibration set. The idea of CQ is identical to the concept of vector quantization (VQ), where groups of values are jointly quantized into codebook indices. Utilizing the centroids as codebooks, CQ achieves state-of-the-art performance in diverse tasks, proving the effectiveness of VQ in KV cache quantization.
In our study, we observe that CQ fails to generalize in out-of-distribution (OOD) scenarios where the input distribution is different from that of the calibration set (in-distribution) (yuan2023benchmarking, ). For example, Figure 1(a) shows that while CQ excels in WikiText-2, it performs much worse in C4, since it is calibrated using only a few samples from WikiText-2 dataset. Moreover, Figure 1(b) demonstrates the distribution of key and value can vary greatly by the input distribution. This implies that learning centroids from the small set of data is very risky when applying them to other datasets. To this end, we propose NSNQuant, a calibration-free vector quantization (VQ) method that generalizes well to a wide range of datasets. Our contributions are summarized as follows:
@@
  • To address this limitation, we propose NSNQuant, a calibration-free VQ method for KV cache. NSNQuant effectively matches the key and value distribution with the standard normal distribution through the Normalize-Shift-Normalize (NSN) process and the Hadamard transform, as shown in Figure 3. Since the approximate distribution of key-value is known prior to the inference, an effective codebook for VQ can be built without any external data.
  • We conduct comprehensive experiments and analysis to show the effectiveness of NSNQuant across different tasks and models. The results with the LLaMA (2024llama3.1, ; touvron2023llama, ; grattafiori2024llama, ) and Mistral models (jiang2023mistral7b, ) clearly show that NSNQuant outperforms other baselines in 1-bit and 2-bit quantization. Moreover, we implement efficient CUDA kernels for the low-bit computation, improving throughput and reducing memory usage.
@@ 2 Preliminary
LLM inference has two main stages: prefilling and decoding. In the prefilling stage, all prompt tokens are processed simultaneously by the transformer decoder layers (vaswani2017attention, ). In the decoding stage, new tokens are generated one by one in an autoregressive manner. In both stages, each token only attends to previous tokens due to the causal nature of masked self-attention. To avoid redundant computation, a KV cache stores key-value pairs from previous tokens. It is initialized during prefilling and updated at each decoding step by appending the latest key-value pair. It becomes a primary bottleneck when processing long sequences, as its size linearly increases with the sequence length.
@@
A Hadamard matrix is an orthogonal matrix whose entries all have the same magnitude. Its size can be doubled recursively through Sylvester’s construction, which builds a larger Hadamard matrix by combining copies of a smaller one in a specific pattern. This recursive definition forms the basis of the Walsh–Hadamard transform, allowing matrix–vector multiplication in $O(d\log d)$ time. QuIP# (pmlr-v235-tseng24a, ) adopts the randomized Hadamard transform (RHT) (halko2011finding, ), where the sign of each row and column is flipped independently with probability 1/2. We also employ RHT to compute the theoretical bounds presented in Lemma 1.
Figure 2: Overall structure of attention under NSNQuant. Residuals are omitted for simplicity. We use superscript $k$ and $v$ to mark values associated with key and value, respectively. Since NSN is applied to keys before RoPE, two branches are needed to correctly compute attention scores. See Appendix B for details on attention output computation.
Figure 3: A visual illustration of the effect of NSN on the channel distribution. Our Normalize-Shift-Normalize (NSN) process aligns the channel distribution with the standard normal distribution, when used with the Hadamard transform.
@@ 3.1 Motivation
As shown in Figure 1(a), we observe several cases where the existing VQ method, coupled quantization (CQ) (zhang2024kv, ), suffers from severe performance degradation when tested on OOD datasets. We attribute this to the distribution mismatch between datasets, as visualized in Figure 1(b). We observe that this mismatch can cause severe errors in certain datasets. A notable example of such errors is the large quantization error observed in punctuation tokens in LLaMA3-8B and LLaMA3.1-8B.
WikiText-2, a calibration dataset for CQ, only contains " , " (with whitespace) tokens, while the C4 dataset includes "," (without whitespace) tokens. This mismatch leads to large errors for the keys corresponding to the punctuation token "," in the first layer of LLaMA3 models, since the centroids of CQ are obtained only from WikiText-2. This leads to significant distortion in attention weights because in certain heads, the ’,’ token accounts for over 90% of the attention weights. By preserving the keys corresponding to these tokens in the first layer, the perplexity of CQ-4c9b on C4 is improved from 13.97 to 9.15 in LLaMA3-8B, and from 12.24 to 9.16 in LLaMA3.1-8B, closely matching the results of NSNQuant.
@@ 3.2 Normalize-Shift-Normalize (NSN)
Table 1: Average channel-wise KL divergence with the standard normal distribution measured on key and value of LLaMA2-7B on WikiText-2 with sequence length of 4096. KL divergence is computed between the binned empirical distribution and the standard normal CDF. We used 4096 random data for Oracle, to match the sequence length.
@@
Let $v\in\mathbb{R}^{l\times d}$ be the tensor where $l$ is a sequence length and $d$ is a hidden dimension per head. In the first Normalize step, each token is normalized to have a norm of $\sqrt{d}$, where $d$ is a token dimension per head. This prevents outlier tokens (chen2024prefixquant, ) from dominating the next steps and incurring large errors in tokens with small magnitude. In the Shift step, channel-wise mean is calculated and subtracted, so that the resulting distribution is zero-centered. Finally, in the second Normalization step, each token is again normalized to have a norm $\sqrt{d}$. The entire process is formulated as follows:
@@
For the randomized Hadamard transform $Y=\operatorname{RHT}(X)$, $10<\alpha<$ 1 and $i\in\{1,2,\dots,d\}$,
[…]
where $\beta_{\alpha}:=\dfrac{1}{d}\sqrt{\dfrac{\ln(2/\alpha)}{c}}$ and $c$ is a constant from the Hanson-Wright inequality (rudelson2013hanson, ).
@@
To implement the second step (Shift) in the decoding stage, we bring the idea of residual from KIVI (liu2024kivi, ). Following KIVI, we split KV cache into two parts: one part with quantized KV cache, and the other part with full-precision KV cache (residual). If the size of the residual reaches its maximum capacity, the KV cache in the residual is flushed, quantized, and appended to the quantized part. We introduce a hyperparameter called residual size to control the size of the residual. To ensure consistency, we also apply NSN in residual-size chunks during the prefilling stage. In our experiments, we set the residual size to 64.
@@ 3.3 Scale adjustment
Let $v\in\mathrm{R}^{d}$ be a token vector which is processed by the NSN and Hadamard transform. It is then divided into 8-dimensional sub-vectors and quantized using the codebook, following the objective in Equation 1. Let $v_{Q}$ be the vector restored by looking up codebook. i.e., $v_{Q}=\mathbb{C}[\operatorname{VQ}(v)]$. We find that rather than using $v_{Q}$ as-is, scaling $v_{Q}$ adaptively improves performance. Specifically, we find that it is beneficial to scale $v_{Q}$ as follows, which is identical to scaling $s_{2}$ since $v_{Q}$ is multiplied by $s_{2}$ when restoring:
$v_{Q}\leftarrow\frac{\|v\|_{2}^{2}}{v\cdot v_{Q}}v_{Q}\quad\text{(i.e., $s_{2}\leftarrow\frac{\|v\|_{2}^{2}}{v\cdot v_{Q}}s_{2}$).}$.
This is a scaling strategy which makes $v_{Q}-v$ orthogonal to $v$. In other words, it preserves components parallel to $v$, while allowing some orthogonal errors. Interpreting $o$ as local context and $v$ as a distinctive token feature, this strategy can be interpreted as making each token distinctive, which is essential for KV cache considering its selective property. We provide comparison for different scaling strategies in Appendix C.3.
@@ 3.4 Codebook tuning
We construct a single global codebook for compressing 8-dimensional vectors using integer indices, following QuIP# (pmlr-v235-tseng24a, ). NSNQuant-2b uses 8 bits for signs and 8 for codebook indices, while NSNQuant-1b uses only 8 bits for indices. A simple baseline can be built via K-Means on standard normal data, but its local optimality limits performance. We improve this by fine-tuning on synthetic standard normal data (torch.randn) to minimize cosine distance between original and quantized vectors, since the error of scale adjustment depends only on the angle between them. As the lookup is non-differentiable, gradients are propagated only through post-lookup operations. The PyTorch implementation for this process runs less than 5 minutes on an RTX 3090, unlike calibration process of CQ or KVQuant which requires backpropagation through model weights. While E8P (pmlr-v235-tseng24a, ) is also a competitive 2-bit baseline, it is hard to apply to the 1-bit scenario, and requires expensive comparisons with over 2000 entries. We provide comparison results in Table 9.
@@ 3.5 Double quantization
To further reduce the memory overhead, we employ double quantization (DQ) proposed in QLoRA (dettmers2023qlora, ), which quantizes parameters used for the quantization. Specifically, we quantize $o$ and $s_{1}$ with a group size of 32 and residual size, respectively, using 4-bit round-to-nearest (RTN) quantization. DQ significantly reduces the average bit width. As a result, NSNQuant costs only additional 0.23 bits on average for the NSN process, when residual size is set to 64. Moreover, to reduce the amount of shared memory required for the codebook, we also apply 4-bit quantization to codebook entries. As shown in Table 11, DQ barely affects the effectiveness of NSNQuant.
@@ 3.6 Efficient kernel implementation
We implement the CUDA kernels for efficient execution of NSNQuant-2b and NSNQuant-1b. For codebook matching, we compute and manage distances on-the-fly in streaming multiprocessors (SMs), while loading codebook tiles into shared memory. For dequantization and matrix-vector multiplications, our kernel loads the codebook into shared memory to minimize the accesses to DRAM. We implement two different kernels for $QK^{T}$ and $WV$ ($W$: attention weights) since they use different axis for reduction. We also fuse the computation of both the quantized and the residual parts to maximize GPU utilization even in the small-batch scenarios.
@@ 4.1 Experimental setup
First, we evaluate perplexity (PPL) on WikiText-2 and C4 dataset to show quantization error in language modeling. Second, we report evaluation results on LongBench (bai-etal-2024-longbench, ). We choose the task subset from LongBench following KIVI (liu2024kivi, ): Qasper for a single-document QA task; QMSum and MultiNews for summarization tasks; TREC, TriviaQA and SAMSum for few-shot learning tasks; LCC and RepoBench-P for code completion tasks. Lastly, we run evaluations on GSM8K (cobbe2021training, ), HumanEval (chen2021codex, ), CoQA (reddy2019coqa, ), and MMLU (hendrycks2020measuring, ) using LM-Eval framework (eval-harness, ) to test generation capability in more diverse generation scenarios.
@@
We compare NSNQuant against four baselines: KIVI (liu2024kivi, ), KIVI + Hadamard, KVQuant (hooper2024kvquant, ), and CQ (zhang2024kv, ). While some of these methods provide multiple versions with similar average bits (e.g., CQ-4c8b and CQ-4c9b), we adopt the stronger variants to better demonstrate the superiority of our approach. Simpler baselines like INT2 are excluded, as prior studies (hooper2024kvquant, ; zhang2024kv, ) have shown them to be ineffective. For NSNQuant, we fix both residual size to 64. For KIVI, we use a group size of 64 for keys and 128 for tokens, to maintain consistency with our residual policy. Detailed explanations for each baseline are provided in Appendix E.3. Note that since CQ does not provide their official implementation, we reproduce it based on the official implementation of KVQuant. The result of CQ is slightly worse than the one reported in the original paper.
@@
Since maintaining full-precision cache largely affects generation quality (liu2024kivi, ; zhang2024kv, ), we unify several settings to focus only on quantization quality: all methods adopt NSNQuant’s residual policy, which buffers full-precision caches until the residual is full. We fix the residual size to 64 in all experiments. Attention sink-aware quantization is removed from KVQuant, as it can be easily applied to other methods as well. Since some methods (e.g., KIVI) use full-precision cache in the prefilling stage, while others (e.g., CQ) use quantized cache, we standardize this by using full-precision cache in the prefilling stage for all methods—except in perplexity evaluations, where key and value quantization is necessary to measure forward-pass quality.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Large language models (LLMs) have been widely adopted across various domains due to their strong generalization capabilities (achiam2023gpt). Recently, with the emerging trend of using LLMs to solve complex problems, their use has expanded into long-context scenarios such as long-context reasoning (wei2022chain; guo2025deepseek) and retrieval-augmented generation (RAG) (lewis2020retrieval). However, when processing long sequences, LLM inference requires a significant amount of memory and is memory-bound (kim2023squeezellm; hooper2024kvquant).
One of the major causes is the large size of key-value (KV) cache, which linearly increases with the sequence length. To alleviate this issue, many studies have proposed approaches to compressing KV cache effectively, such as eviction (zhang2023h2o; li2024snapkv; cai2024pyramidkv; qin2025cake) and low-rank approximation (changpalu; lin2024matryoshkakv). Among them, quantization has been one of the most widely adopted approaches. Previous studies (liu2024kivi; hooper2024kvquant) analyze how outliers emerge in the KV cache and propose to quantize key and value along the channel dimension and token dimension, respectively. Recently, Coupled Quantization (CQ) (zhang2024kv) proposed to quantize multiple channels together, using centroids obtained from the calibration set. The idea of CQ is identical to the concept of vector quantization (VQ), where groups of values are jointly quantized into codebook indices. Utilizing the centroids as codebooks, CQ achieves state-of-the-art performance in diverse tasks, proving the effectiveness of VQ in KV cache quantization.
In our study, we observe that CQ fails to generalize in out-of-distribution (OOD) scenarios where the input distribution is different from that of the calibration set (in-distribution) (yuan2023benchmarking). For example, Figure 1(a) shows that while CQ excels in WikiText-2, it performs much worse in C4, since it is calibrated using only a few samples from WikiText-2 dataset. Moreover, Figure 1(b) demonstrates the distribution of key and value varies greatly with the input distribution. This implies that learning centroids from the small set of data is very risky when applying them to other datasets. To this end, we propose NSNQuant, a calibration-free vector quantization (VQ) method that generalizes well to a wide range of datasets. Our contributions are summarized as follows:
@@
  • To address this limitation, we propose NSNQuant, a calibration-free VQ method for KV cache. NSNQuant effectively matches the key and value distribution with the standard normal distribution through the Normalize-Shift-Normalize (NSN) process and the Hadamard transform, as shown in Figure 3. Since the approximate distribution of key-value is known prior to inference, an effective codebook for VQ can be built without any external data.
  • We conduct comprehensive experiments and analysis to show the effectiveness of NSNQuant across different tasks and models. The results with the LLaMA (2024llama3.1; touvron2023llama; grattafiori2024llama) and Mistral models (jiang2023mistral7b) clearly show that NSNQuant outperforms other baselines in 1-bit and 2-bit quantization. Moreover, we implement efficient CUDA kernels for the low-bit computation, improving throughput and reducing memory usage.
@@ 2 Preliminary
LLM inference has two main stages: prefilling and decoding. In the prefilling stage, all prompt tokens are processed simultaneously by the transformer decoder layers (vaswani2017attention). In the decoding stage, new tokens are generated one by one in an autoregressive manner. In both stages, each token only attends to previous tokens due to the causal nature of masked self-attention. To avoid redundant computation, a KV cache stores key-value pairs from previous tokens. It is initialized during prefilling and updated at each decoding step by appending the latest key-value pair. It becomes a primary bottleneck when processing long sequences, as its size linearly increases with the sequence length.
@@
A Hadamard matrix is an orthogonal matrix whose entries all have the same magnitude. Its size can be doubled recursively through Sylvester’s construction, which builds a larger Hadamard matrix by combining copies of a smaller one in a specific pattern. This recursive definition forms the basis of the Walsh–Hadamard transform, allowing matrix–vector multiplication in $O(d\log d)$ time. QuIP# (pmlr-v235-tseng24a) adopts the randomized Hadamard transform (RHT) (halko2011finding), where the sign of each row and column is flipped independently with probability 1/2. We also employ RHT to compute the theoretical bounds presented in Lemma 1.
Figure 2: Overall structure of attention under NSNQuant. Residuals are omitted for simplicity. We use superscript $k$ and $v$ to mark values associated with key and value, respectively. Since NSN is applied to keys before RoPE, two branches are needed to correctly compute attention scores. Details of the attention computation shown in the figure are provided in Appendix B.
Figure 3: A visual illustration of the effect of NSN on the per-channel value distribution. Our Normalize-Shift-Normalize (NSN) process aligns the distribution with the standard normal distribution, when used with the Hadamard transform.
@@ 3.1 Motivation
As shown in Figure 1(a), we observe several cases where the existing VQ method, coupled quantization (CQ) (zhang2024kv), suffers from severe performance degradation when tested on OOD datasets. We attribute this to the distribution mismatch between datasets, as visualized in Figure 1(b). We observe that this mismatch can cause severe errors in certain datasets. A notable example of such errors is the large quantization error observed in punctuation tokens in LLaMA3-8B and LLaMA3.1-8B.
WikiText-2, a calibration dataset for CQ, only contains " , " (with whitespace) tokens, while the C4 dataset includes "," (without whitespace) tokens. This mismatch leads to large errors for the keys corresponding to the punctuation token "," in the first layer of LLaMA3 models, since the centroids of CQ are obtained only from WikiText-2. This leads to significant distortion in attention weights because in certain heads, the "," token accounts for over 90% of the attention weights. By preserving the keys corresponding to these tokens in the first layer, the perplexity of CQ-4c9b on C4 is improved from 13.97 to 9.15 in LLaMA3-8B, and from 12.24 to 9.16 in LLaMA3.1-8B, closely matching the results of NSNQuant.
@@ 3.2 Normalize-Shift-Normalize (NSN)
Table 1: Average channel-wise KL divergence with the standard normal distribution measured on key and value of LLaMA2-7B on WikiText-2 with sequence length of 4096. KL divergence is computed between the binned empirical distribution and the standard normal CDF. We used 4096 random samples for Oracle, to match the sequence length.
@@
Let $v\in\mathbb{R}^{l\times d}$ be the tensor where $l$ is a sequence length and $d$ is a hidden dimension per head. In the first Normalize step, each token is normalized to have a norm of $\sqrt{d}$, where $d$ is a token dimension per head. This prevents outlier tokens (chen2024prefixquant) from dominating the next steps and incurring large errors in tokens with small magnitude. In the Shift step, channel-wise mean is calculated and subtracted, so that the resulting distribution is zero-centered. Finally, in the second Normalize step, each token is again normalized to have a norm $\sqrt{d}$. The entire process is formulated as follows:
@@
For the randomized Hadamard transform $Y=\operatorname{RHT}(X)$, $0<\alpha<1$ and $i\in\{1,2,\dots,d\}$,
$\operatorname{P}\Bigl(\operatorname{Var}(Y_{i})\!\in\!\bigl[\,1-\varepsilon-\Gamma\,\beta_{\alpha},\;1+\Gamma\,\beta_{\alpha}\bigr]\Bigr)\;\geq\;1-\alpha$
where $\beta_{\alpha}:=\dfrac{1}{d}\sqrt{\dfrac{\ln(2/\alpha)}{c}}\$and $c$ is a constant from the Hanson-Wright inequality (rudelson2013hanson).
@@
To implement the second step (Shift) in the decoding stage, we bring the idea of residual from KIVI (liu2024kivi). Following KIVI, we split KV cache into two parts: one part with quantized KV cache, and the other part with full-precision KV cache (residual). If the size of the residual reaches its maximum capacity, the KV cache in the residual is flushed, quantized, and appended to the quantized part. We introduce a hyperparameter called residual size to control its size. To ensure consistency, we also apply NSN in residual-size chunks during the prefilling stage. In our experiments, we set the residual size to 64.
@@ 3.3 Scale adjustment
Let $v\in\mathrm{R}^{d}$ be a token vector that is processed by the NSN and Hadamard transform. It is then divided into 8-dimensional sub-vectors and quantized using the codebook, following the objective in Equation 1. Let $v_{Q}$ be the vector restored by looking up the codebook. i.e., $v_{Q}=\mathbb{C}[\operatorname{VQ}(v)]$. We find that rather than using $v_{Q}$ as-is, scaling $v_{Q}$ adaptively improves performance. Specifically, we find that it is beneficial to scale $v_{Q}$ as follows, which is identical to scaling $s_{2}$ since $v_{Q}$ is multiplied by $s_{2}$ when restoring:
$v_{Q}\leftarrow\frac{\|v\|_{2}^{2}}{v\cdot v_{Q}}v_{Q}\quad\text{(i.e., $s_{2}\leftarrow\frac{\|v\|_{2}^{2}}{v\cdot v_{Q}}s_{2}$).}$
This is a scaling strategy which makes $v_{Q}-v$ orthogonal to $v$. In other words, it preserves components parallel to $v$, while allowing some orthogonal errors. Interpreting $o$ as local context and $v$ as a distinctive token feature, this strategy can be interpreted as making each token distinctive, which is essential for KV cache considering its selective property. We provide a comparison of different scaling strategies in Appendix C.3.
@@ 3.4 Codebook tuning
We construct a single global codebook for compressing 8-dimensional vectors using integer indices, following QuIP# (pmlr-v235-tseng24a). NSNQuant-2b uses 8 bits for signs and 8 bits for codebook indices, while NSNQuant-1b uses only 8 bits for indices. A simple baseline can be built via K-Means on standard normal data, but its local optimality limits performance. We improve this by fine-tuning on synthetic standard normal data (torch.randn) to minimize cosine distance between original and quantized vectors, since the error of scale adjustment depends only on the angle between them. As the lookup is non-differentiable, gradients are propagated only through post-lookup operations. The PyTorch implementation for this process runs in less than 5 minutes on an RTX 3090, unlike the calibration process of CQ or KVQuant which requires backpropagation through model weights. While E8P (pmlr-v235-tseng24a) is also a competitive 2-bit baseline, it is hard to apply to the 1-bit scenario, and requires expensive comparisons with over 2000 entries. We provide comparison results in Table 10.
@@ 3.5 Double quantization
To further reduce the memory overhead, we employ double quantization (DQ) proposed in QLoRA (dettmers2023qlora), which quantizes parameters used for the quantization. Specifically, we quantize $o$ and $s_{1}$ with a group size of 32 and residual size, respectively, using 4-bit round-to-nearest (RTN) quantization. DQ significantly reduces the average bit width. As a result, NSNQuant costs only additional 0.23 bits on average for the NSN process, when residual size is set to 64. Moreover, to reduce the amount of shared memory required for the codebook, we also apply 4-bit quantization to codebook entries. As shown in Table 12, DQ barely affects the effectiveness of NSNQuant.
@@ 3.6 Efficient kernel implementation
We implement the CUDA kernels for efficient execution of NSNQuant-2b and NSNQuant-1b. For codebook matching, we compute and manage distances on-the-fly in streaming multiprocessors (SMs), while loading codebook tiles into shared memory. For dequantization and matrix-vector multiplications, our kernel loads the codebook into shared memory to minimize access to DRAM. We implement two different matrix-vector multiplication kernels for $qK^{T}$ and WvWv ($W$: attention weights) since they use different axes for reduction. We also fuse the computation of both the quantized and the residual parts to maximize GPU utilization even in the small-batch scenarios.
@@ 4.1 Experimental setup
First, we evaluate perplexity (PPL) on WikiText-2 and C4 dataset to show quantization error in language modeling. Second, we report evaluation results on LongBench (bai-etal-2024-longbench). We choose the task subset from LongBench following KIVI (liu2024kivi): Qasper for a single-document QA task; QMSum and MultiNews for summarization tasks; TREC, TriviaQA and SAMSum for few-shot learning tasks; LCC and RepoBench-P for code completion tasks. Lastly, we run evaluations on GSM8K (cobbe2021training), HumanEval (chen2021codex), CoQA (reddy2019coqa), and MMLU (hendrycks2020measuring) using LM-Eval framework (eval-harness) to test generation capability in more diverse generation scenarios.
@@
We compare NSNQuant against four baselines: KIVI (liu2024kivi), KIVI + Hadamard, KVQuant (hooper2024kvquant), and CQ (zhang2024kv). While some of these methods provide multiple versions with similar average bits (e.g., CQ-4c8b and CQ-4c9b), we adopt the stronger variants to better demonstrate the superiority of our approach. Simpler baselines like INT2 are excluded, as prior studies (hooper2024kvquant; zhang2024kv) have shown them to be ineffective. For NSNQuant, we fix the residual size to 64. For KIVI, we use a group size of 64 for keys and 128 for tokens, to maintain consistency with our residual policy. Detailed explanations for each baseline are provided in Appendix E.3. Note that since CQ does not provide their official implementation, we reproduce it based on the official implementation of KVQuant. The result of CQ is slightly worse than the one reported in the original paper.
@@
Since maintaining full-precision cache largely affects generation quality (liu2024kivi; zhang2024kv), we unify several settings to focus only on quantization quality: all methods adopt NSNQuant’s residual policy, which buffers full-precision caches until the residual is full. We fix the residual size to 64 in all experiments. Attention sink-aware quantization is removed from KVQuant, as it can be easily applied to other methods as well. Since some methods (e.g., KIVI) use full-precision cache in the prefilling stage, while others (e.g., CQ) use quantized cache, we standardize this by using full-precision cache in the prefilling stage for all methods—except in perplexity evaluations, where key and value quantization is necessary to measure forward-pass quality.